OpenStudy (anonymous):

Which of the following is the slope between the points (-1, -5) and (7, -3)?

OpenStudy (anonymous):

2/8

OpenStudy (saifoo.khan):

\[Slope = \frac{Rise}{Run}\]

OpenStudy (anonymous):

slope= y2-y1 / x2-x1

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (anonymous):

I got 1/4 - it's the same thing right?

OpenStudy (saifoo.khan):

how u got it?

OpenStudy (anonymous):

rise over run ^ - I know how to do it, I'm just checking my answers.

OpenStudy (saifoo.khan):

Solve it.

OpenStudy (anonymous):

-3 saif

OpenStudy (saifoo.khan):

oh Right.

OpenStudy (saifoo.khan):

Rise / Run (-1, -5) and (7, -3) (x , y) and (x,y) \[\frac{-3-(-5)}{7-(-1)}\]

OpenStudy (anonymous):

ok , wt do u mean by rise & run??

OpenStudy (anonymous):

I know it equals to 2/8, but that's not an option, that why I reduced it to get 1/4. lol..

OpenStudy (anonymous):

yup 2/8 = 1/4

OpenStudy (saifoo.khan):

@ruba, try plotting it on graph. it is just a method taught to new learners so they wont forget it.

OpenStudy (anonymous):

ok so wt z rise nd run

OpenStudy (saifoo.khan):

Rise means change in y axis over change in x axis

OpenStudy (anonymous):

thx